# 22-Split of Form Factors via BCFW Recursion Relation

### v1

@@ 1 Introduction
The structure of this note is as follows. In Section 2, we identify the hidden zeros and a new zero relevant to the form factors of ${\cal O}=\tfrac{1}{2}\mathrm{Tr}[(\partial\phi)^{2}]+\mathrm{Tr}(\phi^{3})$ and ${\cal O}=\mathrm{Tr}(F^{2})$. In Section 3, we employ these results within an inductive proof based on BCFW recursion relation, thereby establishing the $2$-split factorization for arbitrary multiplicity. We conclude in Section 4 with a brief summary and outlook.
@@ 2 Hidden zeros
Let us consider the operator ${\cal O}={\cal L}_{\phi}\equiv\frac{1}{2}\rm Tr\left((\partial\phi)^{2}\right)+\rm Tr(\phi^{3})$ with external particles taken to be scalars in the $\mathrm{Tr}(\phi^{3})$ theory. We select three distinguished particles $i,j,k$ and partition the remaining external legs into two disjoint sets $A$ and $B$, such that $A\cup B=\{1,\ldots,n\}\setminus\{i,j,k\}$. We then impose the constraint
@@
where $s_{a,b}=2p_{a}\cdot p_{b}$ and $s_{a,q}=2p_{a}\cdot q$. Under these kinematic conditions, the form factor $F^{{\cal L}_{\phi}}_{n}(1,\ldots,n;q)$ exhibits a $2$-split factorization:
$F^{{\cal L}_{\phi}}_{n}(i,A,j,B(k);q)\to{\cal A}^{\phi^{3}}_{|A|+3}(i,A,j;\kappa)F^{{\cal L}_{\phi}}_{|B|+3}(j,B,i;\kappa^{\prime},q)\,$, (3)
where $B(k)\equiv B\cup\{k\}$ preserves the ordering [^1] For instance, $B=\{1,2,4\},k=3$, the $B(k)=\{1,2,3,4\}$.. We use the semicolon to single out the off-shell leg, which is not written according to the ordering. $\kappa,\kappa^{\prime}$ take the position of that of the leg kk [^2] For example, $k=n$, then $(i,A,j;\kappa)=(i,A,j,\kappa)$ and $(j,B,j;\kappa^{\prime})=(j,\ldots,n-1,\kappa^{\prime},1,\ldots,i)$..
@@
Assuming (6) holds for form factors with fewer than $n$ points, we can establish it for $F^{{\cal L}_{\phi}}_{n}$ using the BCFW recursion relation. First, we must confirm that the form factor has no boundary contribution under a BCFW shift:
$p_{i}\to\hat{p}_{i}\to p_{i}+zr,~~~~~~p_{j}\to\hat{p}_{j}\to p_{j}+zr$ (8)
with $r^{2}=r\cdot p_{i}=r\cdot p_{j}=0$. The primary difference between this form factor and a pure amplitude arises from the kinematic term $\frac{1}{2}\rm Tr\left((\partial\phi)^{2}\right)$ in the operator ${\cal L}_{\phi}$. This introduces a non-trivial numerator $p\cdot q$, but it also provides an additional propagator. The net scaling at large $z$ is therefore favorable, allowing the BCFW recursion relation to be applied in the form factor Brandhuber:2010ad :
@@
In the kinematic limit where $s_{a,q}=0$ for all $a\in\alpha$, every term in the BCFW sum involves a lower-point form factor $F^{{\cal L}_{\phi}}_{L}$ or $F^{{\cal L}_{\phi}}_{R}$ evaluated on a subset of these legs. By the induction hypothesis, these lower-point form factors vanish. Consequently, the entire right-hand side is zero.
This completes the proof, confirming that (6) is valid for all $n\geqslant 3$.
Similarly, the Yang-Mills (YM) form factor for the operator ${\cal O}=\rm Tr(F^{2})$ possesses hidden zeros. The YM form factor also exhibits a $2$-split factorization on a different kinematic locus:
@@ 3.1 ${\cal O}=\frac{1}{2}\rm Tr\left((\partial\phi)^{2}\right)+\rm Tr(\phi^{3})$
Summing the contributions from these four non-vanishing diagrams in (17), (3.1), (3.1), and (3.1) precisely reconstructs the BCFW expansion for the right-hand side of the $2$-split formula, (3). The first 2 diagrams build the right-hand current, while the last 2 diagrams build the right-hand form factor.
This completes the inductive proof for the $2$-split of the form factor $F^{{\cal L}_{\phi}}_{n}$ (3).

### v2

@@ 1 Introduction
The structure of this paper is as follows. In Section 2, we present a brief review of the $2$-split behavior. In Section 3, we identify the hidden zeros and a new zero relevant to the form factors of ${\cal O}=\tfrac{1}{2}{\rm Tr}[(\partial\phi)^{2}]+{\rm Tr}(\phi^{3})$ and ${\cal O}={\rm Tr}(F^{2})$. In Section 4, we employ these results within an inductive proof based on BCFW recursion relation, thereby establishing the $2$-split factorization for arbitrary multiplicity. We conclude in Section 5 with a brief summary and outlook.
2 Review
Let us first briefly review the so-called $2$-split behavior discovered in Cao:2024gln ; Cao:2024qpp , where certain amplitudes factorize into two currents under specific kinematic constraints. Consider the ${\rm Tr}(\phi^{3})$ theory as an example. Select three distinguished particles $i,j,k$ and partition the remaining external legs into two disjoint sets $A$ and $B$ such that $A\cup B=\{1,\ldots,n\}\setminus\{i,j,k\}$. Impose the constraint
$s_{a,b}=0\,,\quad\forall a\in A,b\in B\,$, (2)
where the Mandelstam invariants are $s_{a,\ldots,m}=(p_{a}+\ldots+p_{m})^{2}$. Under this condition, the amplitude ${\cal A}^{\phi^{3}}_{n}(i,A,j,B(k))$ exhibits a $2$-split factorization:
${\cal A}^{\phi^{3}}_{n}(i,A,j,B(k))\to{\cal J}^{\phi^{3}}_{|A|+3}(i,A,j;\kappa){\cal J}^{\phi^{3}}_{|B|+3}(j,B,i;\kappa^{\prime})\,$, (3)
where $B(k)\equiv B\cup\{k\}$ preserves the ordering [^1] For instance, $B=\{1,2,4\},k=3$, the $B(k)=\{1,2,3,4\}$.. The semicolon indicates the off-shell leg, not shoen in the ordering, and $\kappa,\kappa^{\prime}$ take the position of the leg kk [^2] For example, if $k=n$, then $(i,A,j;\kappa)=(i,A,j,\kappa)$ and $(j,B,j;\kappa^{\prime})=(j,\ldots,n-1,\kappa^{\prime},1,\ldots,i)$..
As an illustration, the $5$-point amplitude is
${\cal A}^{\phi^{3}}(2,3,4,5,1)=\frac{1}{s_{1,2}}\frac{1}{s_{1,2,3}}+\frac{1}{s_{2,3}}\frac{1}{s_{1,2,3}}+\frac{1}{s_{2,3}}\frac{1}{s_{2,3,4}}+\frac{1}{s_{1,2}}\frac{1}{s_{3,4}}+\frac{1}{s_{3,4}}\frac{1}{s_{2,3,4}}\,.$ (4)
Setting $i=2,j=4,k=5$, and applying $s_{a,b}=s_{3,1}=0$, the amplitude factorizes as
${\cal A}^{\phi^{3}}(2,3,4,5,1)\to$ $\left(\frac{1}{s_{2,3}}+\frac{1}{s_{3,4}}\right)\,\left(\frac{1}{s_{1,2}}+\frac{1}{s_{1,5}}\right)$
$\equiv$ ${\cal A}^{\phi^{3}}(2,3,4;\kappa){\cal A}^{\phi^{3}}(4,1,2;\kappa^{\prime})\,.$ (5)
For Yang-Mills (YM) amplitudes, additional polarization constraints are required:
$\epsilon_{a}\cdot\epsilon_{b^{\prime}}=p_{a}\cdot\epsilon_{b^{\prime}}=\epsilon_{a}\cdot p_{b}=0\,$, (6)
where $b^{\prime}\in B^{\prime}=B\cup\{i,j,k\}$. Under the constraints (2) and (6), the YM amplitude factorizes as
${\cal A}^{\rm YM}(i,A,j,B(k))\to{\cal J}^{\rm mixed}_{|A|+3}(i^{\phi},A,j^{\phi};\kappa^{\phi})\,{\cal J}^{\rm YM}_{|B|+3}(j,B,i;\kappa^{\prime})_{\mu}\epsilon_{k}^{\mu}\,$, (7)
where the pure gluon current contracts with polarization $\epsilon_{k}$. As a simple example, consider the $4$-point Yang–Mills amplitude with $i=1$, $j=3$, $k=4$, and A=2A={2}. Under the $2$-split kinematic and polarization constraints, the amplitude factorizes as
${\cal A}^{\rm YM}(1,2,3,4)\to$ $\left(\frac{p_{1}\cdot\epsilon_{2}}{s_{12}}-\frac{p_{3}\cdot\epsilon_{2}}{s_{2,3}}\right)\left(\epsilon_{1}\cdot\epsilon_{3}p_{3,\mu}+p_{1}\cdot\epsilon_{3}\epsilon_{1,\mu}-p_{3}\cdot\epsilon_{1}\epsilon_{3,\mu}\right)\epsilon_{4}^{\mu}$
$\equiv{\cal J}^{\rm mixed}(1^{\phi},2,3^{\phi};\kappa^{\phi}){\cal J}_{\mu}^{\rm YM}(1,3;\kappa^{\prime})\epsilon_{4}^{\mu}\,.$ (8)
More examples and effective theory amplitudes can be found in Cao:2024gln ; Cao:2024qpp .
@@ 3 Hidden zeros
We employ the BCFW recursion relation to prove the $2$-split behavior of form factors. A key question is whether the shifted form factor vanishes as $z\to\infty$ under the BCFW shift,
$p_{i}\to\hat{p}_{i}=p_{i}+zr,~~~~~~p_{j}\to\hat{p}_{j}=p_{j}+zr$ (9)
with $r^{2}=r\cdot p_{i}=r\cdot p_{j}=0$.
We decompose the operator $\cal O$} into two parts: a kinematic term ${\cal O}_{\rm kin}$ and an interaction term ${\cal O}_{\rm int}$. The insertion of ${\cal O}_{\rm int}$ makes the form factor behave analogously to the corresponding amplitude, while the essential difference between a form factor and a pure amplitude arises from ${\cal O}_{\rm kin}$. This term introduces a non-trivial numerator but also provides an additional propagator. Consequently, the relation between a form factor and the corresponding amplitude takes the form
$F_{n}^{\cal O}=F^{{\cal O}_{\rm kin}}_{n}+F^{{\cal O}_{\rm int}}_{n}=\sum_{i}\frac{V^{{\cal O}_{\rm kin}}}{(p_{i}+q)^{2}}{\cal A}_{n}+C{\cal A}_{n}\,$, (10)
where $C$ is a constant, and the summation runs over all external and internal legs at which the operator ${\cal O}_{\rm kin}$ is inserted. Therefore, our scaling analysis reduces to examining the large-$z$ behavior of the product formed by the additional numerator and the propagator arising from the insertion of ${\cal O}_{\rm kin}$.
As an example, consider the operator ${\cal O}={\cal L}_{\phi}\equiv\frac{1}{2}{\rm Tr}\left[(\partial\phi)^{2}\right]+{\rm Tr}(\phi^{3})$. Here ${\cal O}_{\rm kin}=\frac{1}{2}{\rm Tr}\left[(\partial\phi)^{2}\right]$ and ${\cal O}_{\rm int}={\rm Tr}(\phi^{3})$. The $3$-point form factor is given by
$F^{{\cal L}_{\phi}}_{3}(1,2,3;q)=$ $F^{{\cal O}_{\rm kin}}_{3}(1,2,3;q)+F^{{\cal O}_{\rm int}}_{3}(1,2,3;q)$
$=$ $\frac{s_{1,2}+s_{1,3}}{s_{2,3}}+\frac{s_{1,2}+s_{2,3}}{s_{1,3}}+\frac{s_{1,3}+s_{2,3}}{s_{1,2}}+2$
$=$ $\frac{(s_{1,2}+s_{2,3})(s_{1,2}+s_{1,3})(s_{1,3}+s_{2,3})}{(s_{1,2}s_{1,3}s_{2,3})}\,.$ (11)
For the amplitude ${\cal A}^{\phi^{3}}$, since the $i$ and $j$ are non-adjacent, the large $z$ scaling is at least $z^{-1}$. For the YM amplitude ${\cal A}^{\rm YM}$, the scaling is $z^{-2}$ Arkani-Hamed:2008bsc . We now examine whether the product of the additional vertex $V^{\cal O}(m,n;q)$ and the extra propagator alters this scaling and produces a non-vanishing boundary term.
For the operator ${\cal L}_{\phi}$, the extra vertex behave as $V^{{\cal L}_{\phi}}(m,n;q)\propto p_{m}\cdot q\sim z$, while the additional propagator scales as $z^{-1}$ for large $z$. Hence, the overall scaling of $F_{n}^{{\cal L}_{\phi}}$ is $z^{-1}$, identical to ${\cal A}^{\phi^{3}}$. For ${\cal O}=\rm Tr(F^{2})$, the additional vertex is $V_{\mu\nu}^{{\rm Tr}(F^{2})}(m,n;q)=p_{m}\cdot qg_{\mu\nu}-p_{m,\mu}q_{\nu}$, which scales as $z^{2}$ at most when contracted with a momentum $p$ or polarization $\epsilon$. The extra propagator scales as $z^{-1}$, while ${\cal A}^{\rm YM}$ scales as $z^{-2}$. Therefore, the overall scaling of $F_{n}^{{\rm Tr}(F^{2})}$ is again $z^{-1}$. Consequently, for the form factors considered in this paper, the boundary term vanishes.
Analogous to the BCFW recursion relation in Feng:2025ofq , if the form factor possesses hidden zeros, the recursion formula can be greatly simplified. Consider the operator ${\cal L}_{\phi}$ with external particles taken to be scalars in the $\mathrm{Tr}(\phi^{3})$ theory. Analogous to the amplitude, the form factor $F^{{\cal L}_{\phi}}_{n}(1,\ldots,n;q)$ exhibits a $2$-split factorization:
$F^{{\cal L}_{\phi}}_{n}(i,A,j,B(k);q)\to{\cal A}^{\phi^{3}}_{|A|+3}(i,A,j;\kappa)F^{{\cal L}_{\phi}}_{|B|+3}(j,B,i;\kappa^{\prime},q)\,.$ (12)
under these kinematic conditions
@@
where $s_{a,q}=2p_{a}\cdot q$.
@@
Assuming (16) holds for form factors with fewer than $n$ points, we can establish it for $F^{{\cal L}_{\phi}}_{n}$ using the BCFW recursion relation. Under the BCFW shift (9), We apply the BCFW recursion relation in the form factor Brandhuber:2010ad :
@@
In the kinematic limit where $s_{a,q}=0$ for all $a\in\alpha$, every term in the BCFW sum involves a lower-point form factor $F^{{\cal L}_{\phi}}_{L}$ or $F^{{\cal L}_{\phi}}_{R}$ evaluated on a subset of these legs. By the induction hypothesis, these lower-point form factors vanish. Consequently, the entire right-hand side is zero. This completes the proof, confirming that (16) is valid for all $n\geqslant 3$.
Similarly, YM form factor for the operator ${\cal O}=\rm Tr(F^{2})$ possesses hidden zeros. The YM form factor also exhibits a $2$-split factorization on a different kinematic locus:
@@ 4.1 ${\cal O}=\frac{1}{2}\rm Tr\left((\partial\phi)^{2}\right)+\rm Tr(\phi^{3})$
Summing the contributions from these four non-vanishing diagrams in (26), (4.1), (4.1), and (4.1) precisely reconstructs the BCFW expansion for the right-hand side of the $2$-split formula, (12). The first 2 diagrams build the right-hand current, while the last 2 diagrams build the right-hand form factor. This completes the inductive proof for the $2$-split of the form factor $F^{{\cal L}_{\phi}}_{n}$ (12).